# PDE-Transformer: A Continuous Dynamical Systems Approach to Sequence Modeling

### v2

@@ Abstract
We propose PDE-Transformer, a novel sequence‐modeling paradigm that casts the forward pass of a Transformer as the numerical discretization of a continuous reaction–diffusion system derived from a variational energy functional. In our framework, token embeddings evolve under a partial differential equation whose nonlocal integral term models self-attention, local reaction term models feed-forward layers, diffusion term encodes positional smoothing, and a stability control term corresponds to layer normalization. From this unifying perspective, we design an Adaptive PDE Diffusion Layer—an efficient, learnable finite-difference stencil that enforces local smoothness in feature space with linear time complexity and complements self-attention’s global routing. Through a systematic theoretical analysis based on four pillars (stability, diffusion geometry, multi-scale dynamics, and component coupling), we derive principled guidelines for integrating the PDE layer at seven candidate points in the Transformer. Empirically, on the Long Range Arena benchmark, placing the layer immediately after embedding yields a 4.1 pp average accuracy gain over a strong baseline, and an adaptive multi-scale variant delivers further improvements. Our work thus offers a principled, lightweight mechanism to bolster long-range dependency modeling by harmonizing continuous PDE smoothing with discrete self-attention.
@@ 1 Introduction
Since its inception, the Transformer architecture has achieved revolutionary breakthroughs across diverse fields, from natural language processing and computer vision to protein folding, owing to its powerful self-attention mechanism. However, as applications scale from short texts to ultra-long sequences—such as document-level dialogues or whole-genome sequences—the standard Transformer reveals two fundamental bottlenecks. First, its computational and memory complexity, which scales quadratically with sequence length $L$ ($O(L^{2})$), becomes prohibitively expensive. Second, its purely content-driven global interaction mechanism lacks explicit modeling of local geometric structure, hindering its ability to effectively capture long-range dependencies. Although variants like sparse attention and low-rank approximations offer patchwork optimizations, they remain engineering modifications to a discrete computational graph, failing to transcend the limitations of the discrete paradigm or provide a unified theoretical framework for their design.
To overcome this impasse, we propose a return to the first principles of physics, reframing the process of sequence modeling as a continuous variational dynamical system. In this framework, token embeddings do not undergo discrete layered transformations but rather evolve within an energy field governed by three fundamental forces: local diffusion, nonlinear reaction, and nonlocal coupling. We establish a one-to-one correspondence between these forces and the core components of the Transformer: the nonlinear reaction maps to the feed-forward network (FFN), and the nonlocal coupling maps to the self-attention mechanism. This perspective strikingly reveals a structural deficiency in the standard Transformer: the absence of the crucial diffusion term. In physical systems, this term is responsible for penalizing sharp variations and imposing local smoothness, which is key to forming stable, ordered structures.
Building on this core finding, we design a lightweight, plug-and-play Adaptive PDE Diffusion Layer with linear complexity. This module discretizes a reaction-diffusion equation via a learnable finite-difference method, introducing a structured local smoothness inductive bias into the model. It forms a natural complementarity with self-attention: the PDE diffusion layer is responsible for reinforcing local geometric consistency, while self-attention focuses on capturing global content-based associations.
To validate this paradigm, we systematically investigate the impact of integrating the PDE diffusion layer at seven distinct points within the Transformer architecture. Experiments on the challenging Long Range Arena (LRA) benchmark provide compelling evidence for our theory: placing the PDE layer immediately after the initial token embeddings and before the first Transformer block yields the most significant performance gains. This configuration achieves an average accuracy improvement of 4.1 percentage points over a strong baseline, with a multi-scale variant delivering further improvements. This finding uncovers a core mechanism: before global, sparse attention interactions can be effective, the raw semantic space must first undergo local, dense, structured smoothing.
The main contributions of this work are fourfold. First, we introduce a new theoretical paradigm that unifies sequence modeling as a continuous reaction–diffusion dynamical system derived from variational principles, offering a novel physical lens to understand and improve Transformer architectures. Second, we design the Adaptive PDE Diffusion Layer as a plug-and-play, linear-complexity module that significantly enhances the model’s ability to capture local structure at negligible computational cost. Third, we conduct a systematic empirical study to identify the optimal integration point for such a local smoothing mechanism within the Transformer. Finally, we reveal the profound complementarity between explicit local geometric smoothing and global content-based aggregation, providing principled guidance for building more robust and efficient long-sequence models.
The remainder of this paper is organized as follows. Section 2 reviews prior work on efficient Transformers and continuous-time models. Section 3 details our theoretical framework, deriving the PDE-Transformer from first principles. Section 4 presents our comprehensive experimental validation on the LRA benchmark. Finally, Section 5 concludes the paper and discusses future directions.

### v3

# Where to Add PDE Diffusion in Transformers

@@ Abstract
Transformers achieve powerful content-based global routing via self-attention, but they do not explicitly encode a local geometric prior along the sequence axis, making the placement of locality-inducing modules largely empirical in prior hybrid designs. We study a simple, deterministic PDE diffusion layer—one explicit Euler step of 1D heat-equation smoothing implemented by a discrete Neumann Laplacian under a spectral stability budget—and ask a structural question: where should diffusion be inserted relative to attention? Our central claim is that diffusion and attention generally do not commute, so inserting the same local operator before versus after attention can induce qualitatively different behaviors. To formalize this, we develop a three-layer operator-theoretic framework that (i) provides unconditional guarantees for the diffusion subsystem (spectral non-expansiveness and monotone Dirichlet-energy dissipation under $\alpha<1/2$), (ii) derives compositional perturbation bounds linking insertion effects to representation roughness and downstream amplification, and (iii) uses diffusion–attention non-commutativity as a diagnostic for structural “double-mixing” conflicts. Guided by these predictions, we evaluate seven canonical insertion positions on Long Range Arena and find that early diffusion acts as effective pre-regularization, improving average accuracy by +4.1+4.1 points when applied after embedding, whereas post-attention diffusion degrades performance by 2.5-2.5 points, consistent with the predicted conflict. We further show that a multi-scale diffusion variant yields consistent gains under the same global stability budget. Beyond this specific module, our analysis provides a transferable template for reasoning about local–global compositions in sequence models by cleanly separating provable guarantees, compositional bounds, and mechanistic diagnostics.
Transformers, locality, PDE diffusion, operator theory, spectral stability
@@ 1 Introduction
1.1 Motivation
Transformers have become the dominant architecture for sequence modeling, largely due to self-attention’s ability to perform content-based global routing (Vaswani et al., 2017). By enabling interactions between arbitrary token pairs, attention efficiently captures long-range dependencies and global semantic structure. However, standard Transformers lack an explicit local geometric prior along the sequence dimension: they do not directly encourage neighboring token representations to vary smoothly or preserve local coherence. While attention can learn locality implicitly, it is not encoded architecturally and may be brittle under distribution shifts (e.g., different sequence lengths).
This gap has motivated hybrid designs that inject locality through convolutional layers (Wu et al., 2019; Dai et al., 2020), sparse patterns (Beltagy et al., 2020), or other structured inductive biases. Yet a surprisingly basic question remains under-theorized: where should a local operator be placed relative to attention? In practice, placement is often treated as a hyperparameter and tuned by validation performance. In this work, we ask whether placement can be predicted from structure rather than guessed from sweeps: can operator theory explain when local smoothing helps and when it hurts?
Our central observation is simple but consequential: local smoothing and content-based global routing are, in general, non-commuting operations. Let $\mathcal{S}$ denote a local PDE diffusion (smoothing) operator and $\mathcal{A}$ an attention-based routing operator. Typically, $\mathcal{S}\circ\mathcal{A}\neq\mathcal{A}\circ\mathcal{S}$, and the commutator $\mathcal{S},\mathcal{A}]=\mathcal{S}\mathcal{A}-\mathcal{A}\mathcal{S}$ captures this incompatibility. This has direct architectural implications: applying PDE diffusion before attention can pre-regularize a rough representation space and provide a smoother manifold for routing; applying PDE diffusion after attention can instead blur structures just created by non-local mixing, inducing a harmful double-mixing effect.
1.2 Our Approach: A Three-Layer Operator-Theoretic Framework
To formalize these intuitions, we develop a systematic operator-theoretic framework for analyzing interactions between locality and attention in Transformers. The framework is organized into three conceptual layers, each addressing a different aspect of the problem.
Layer I: Intrinsic stability (provable guarantees).
We first study the PDE diffusion operator in isolation. Using a discrete Laplacian with a strict spectral budget constraint ($\alpha<1/2$), we establish unconditional guarantees for the PDE diffusion subsystem alone. Specifically, we show that the operator is spectrally non-expansive, with strict contraction on non-constant modes, and that it induces monotone dissipation of the discrete Dirichlet energy. These results formalize PDE diffusion as a stable geometric low-pass filter that reduces local roughness independently of the Transformer’s non-convex components, thereby providing a mathematically grounded stability anchor. Takeaway: under a simple spectral budget, PDE diffusion is a certified non-expansive local smoother.
Layer II: Compositional analysis (perturbation bounds).
We then analyze how this stable PDE diffusion operator interacts with the rest of the Transformer when composed at different locations. Treating downstream modules—such as LayerNorm, attention, and MLPs—as locally Lipschitz mappings, we derive perturbation propagation bounds showing that the effect of inserting PDE diffusion at a given position depends on two factors: the geometric roughness of the representation at the insertion point, and the sensitivity of subsequent layers to perturbations. This analysis explains why the same PDE diffusion operator can be beneficial when applied early, where representations are rough and regularization is helpful, yet detrimental when applied late, where representations are already smooth and additional smoothing becomes unnecessary or harmful. Takeaway: PDE diffusion helps most where representations are rough and downstream amplification is limited—typically early in the network.
Layer III: Non-commutativity diagnostics (design guidelines).
Finally, we turn to the interaction between PDE diffusion and attention. We formalize their incompatibility through the commutator $\mathcal{S},\mathcal{A}](X)$, which measures the representational discrepancy between pre-attention and post-attention diffusion. Under a local linearization of attention, this conflict scales with the mismatch between the local geometric structure enforced by PDE diffusion and the non-local routing patterns produced by attention. This perspective yields concrete, testable design guidelines: PDE diffusion should be applied early as a form of semantic manifold pre-regularization; post-attention PDE diffusion should be avoided due to structural conflict; and multi-scale PDE diffusion, when constrained by a global spectral budget, can improve robustness by covering multiple frequency bands without sacrificing stability. Takeaway: because PDE diffusion and attention generally do not commute, post-attention smoothing can erase structures created by global routing.
1.3 Contributions
This paper makes four main contributions. First, we develop a three-layer operator-theoretic framework that separates provable intrinsic guarantees from compositional perturbation analysis and architectural design heuristics, clarifying what can and cannot be established rigorously. Second, we derive principled architectural guidelines for inserting locality-inducing modules into Transformers, predicting that early-stage insertion should systematically outperform post-attention insertion due to fundamental operator incompatibility. Third, we provide controlled empirical validation on the Long Range Arena benchmark (Tay et al., 2020b), systematically evaluating seven insertion positions. The results align closely with theory: PDE diffusion after embedding yields a +4.1+4.1 percentage point improvement, while PDE diffusion after attention leads to a 2.5-2.5 point degradation. Finally, beyond the specific PDE diffusion operator studied here, we demonstrate how operator-theoretic reasoning can guide the design of hybrid architectures, offering a transferable methodology applicable to other local–global composition problems, such as convolution–attention or state-space–attention hybrids.
1.4 Roadmap
The remainder of the paper is organized as follows. Section 2 reviews related hybrid architectures and theoretical perspectives. Section 3 introduces the operator-theoretic framework and defines the PDE diffusion layer precisely. Section 4 presents the three-layer theoretical analysis, including intrinsic stability guarantees and compositional perturbation bounds. Section 5 empirically validates the resulting architectural guidelines on Long Range Arena. Section 6 concludes.